A conductor has a conductance of 50μS. What is its resistance?
1
Expert's answer
2020-09-30T12:22:17-0400

solution

conductance(G)="50\mu S"

"=50\times10^{-6}S"


so

resistance(R)="\frac{1}{G}"

"then\\R=\frac{1}{50\times10^{-6}}\\R=2\times10^4\Omega"


so resistance is "2\times 10^4\Omega" .